OpenStudy (anonymous):

(3x-3)^(1/2) + (7x+8)^(1/2)=(17x +13)^(1/2)

OpenStudy (shubhamsrg):

square both sides.. keep (something)^1/2 on 1 side and the rest on the other.. square again you should come up with a quadratic eqn..

OpenStudy (chaise):

Power half indicates square root. ^_^ \[\sqrt{3x-3}+\sqrt{7x+8}=\sqrt{17x+13}\] \[\sqrt{3x-3}^2+\sqrt{7x+8}^2=\sqrt{17x+13}^2\] \[3x-3+7x+8=17x+13\] 10x+5=17x+13 -7x+8 x=-8/7

OpenStudy (shubhamsrg):

@chaise what are you doing!!

OpenStudy (shubhamsrg):

:O

OpenStudy (chaise):

x^(1/2) = sqrt(x) x=sqrt(x)^2. What you do to one side of the equation, make sure you do to the other.

OpenStudy (shubhamsrg):

just check your second step!! note that (a+b)^2 is not a^2 + b^2

OpenStudy (chaise):

square root and power half cancel eachother out. I have no idea what you are on about. My answer is correct, as far as I'm aware.

OpenStudy (shubhamsrg):

(3x-3)^(1/2) + (7x+8)^(1/2)=(17x +13)^(1/2) when you sq both sides you have 3x -3 + 7x + 8 + 2 sqrt [(3x-3)(7x+8)] = 17 x +13 which is quite different from what you wrote..

OpenStudy (shubhamsrg):

-8/7 cant be the ans...its out of domain of sqrt(3x-3) !!

OpenStudy (shubhamsrg):

altough am too lazy solving,,but i now what i say is correct here..

OpenStudy (shubhamsrg):

@chaise hope its clear? i mean i maybe wrong (no chance here :P )..but correct me in that case..

OpenStudy (anonymous):

So, square (3x-3) + (7x+8)^1/2 = 17x+13? Then square again? (3x-3) + (7x+8) = (17x+13)^2?

OpenStudy (shubhamsrg):

@chaise i'll tell you whats happening here.. x=-8/7 will give a complex solution,,i dont think its the required case here and your method yields -8/7 because in sqrt(7x +8) , x=-8/7 makes it 0..so 2sqrt[(3x-3)(7x-8)]=0 ,,thus even your mistake yielded a complex root.. what about the only real solution? does your method provide that? mind you,,its not gonna work every time..

OpenStudy (shubhamsrg):

@jazzylj i'll use this let 3x- 3=a ,, 7x - 8 =b and 17x + 13 =c so you have sqrt(a) + sqrt(b) = sqrt(c) on squaring,,you get a + b + 2sqrt(ab) = c => 2 sqrt(ab) = c-a-b now square both sides.. 4ab = something this is how you'll have to do..